OpenStudy (anonymous):

Find the point-slope form of the equation of the line passing through the points (–1, –3) and (4, 1).

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-1\quad ,&-3)\quad &(4\quad ,&1) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values}\)

OpenStudy (bmorse):

1-(-3)/4-(-1)=4/5 this is your slope so now take either points above to solve equation y-1=4/5(x-4) y=1+4/5x-16/5 y=4/5x-11/5

OpenStudy (anonymous):